# Macroeconomics of Racial Disparities: Discrimination, Labor Market, and Wealth

### v2

@@ 3.1 Preferences
We set the period utility function $u(c)$ to be $\log(c)$. The discount factor, $\beta$, is calibrated to match the quarterly capital-output ratio of 10.26, the value used by a number of studies including Den Haan et al. (2010) and Carroll et al. (2017). The resulting value of $\beta$ is 0.9943, and the corresponding quarterly real interest rate is 1.3%.
@@ 3.2 Production
The worker-firm match produces according to a Cobb-Douglas production function, $f(k)=k^{\alpha_{k}}$. We choose $\alpha_{k}$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015, following Nakajima (2012).
@@ 3.3 Productivity and wealth shocks
We use hourly real wage from the Panel Study of Income Dynamics (PSID) to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process. Our estimation strategy closely follows Setty and Yedid-Levi (2021). We run a standard Mincer wage regression of log wages on the demographic controls, including education, labor market experience, race, marital status, year, and state fixed effects for a sample of men 25 years and older. We then use the obtained residuals to estimate an AR(1) regression by utilizing the panel dimension of the PSID data. The estimated regression is at a biennial frequency, given that the PSID data is available once every two years. We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process. The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency. We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, as equal to 0.1680.555We follow Setty and Yedid-Levi (2021) to have only men in the sample. We also calculate the AR(1) processes, including men and women from PSID. The resulting $\rho_{s}$ is unchanged from the men-only sample (0.9411), and $\sigma_{s}$ is slightly lower (0.1671) than the men-only sample.
Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people having zero wealth. The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0086 for white workers. They capture the empirical moments that around 18% of black workers and 7% white workers have zero wealth (Nakajima, 2021).
@@ 3.4 Unemployment insurance
The unemployment insurance system in our model is characterized by the replacement rate, $h$, maximum insurance payout, $\chi$, and the probability of maintaining the eligibility status, $P_{e}$. We choose the unemployment insurance replacement rate, $h$, to be 0.4 along the lines of Shimer (2005), Mitman and Rabinovich (2015), and others. Following Setty and Yedid-Levi (2021), we calibrate the maximum payout $\chi$ to be 0.8433, which amounts to 48% of the median wage in the model. The eligibility probability $P_{e}$ is chosen to be 0.5385 to generate an average unemployment benefit duration of 26 weeks, as in Mitman and Rabinovich (2015).
@@ 3.5 Labor search and wage bargain
The labor market turnover statistics are obtained from Cajner et al. (2017). The elasticity of the matching function, $\iota$, targets the job-finding rate of the black workers, and our calibrated value of 1.3012 is close to that of 1.25 in Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018). We choose firm-type specific job destruction shocks $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers, respectively. The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively. [^6] All the black workers work only with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of the black workers. On the other hand, white workers work in both $p$ and $np$ firms; hence, $\lambda_{p}$ (0.0268) is smaller than the aggregate separation rate of the white workers (0.0380). The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job-finding rate of white workers, while the posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to target the aggregate labor market tightness, $\theta$, of 1.0, following Wolcott (2021). [^7] Aggregate labor market tightness $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$. We find that the prejudiced firms pay a lot more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated to be 4.2622, compared to $\kappa_{np}$ taking a value of 2.735. The bargaining power of black workers, $\xi_{bl}$, is calibrated to target the average black-white racial wage ratio of 0.75 (Derenoncourt and Montialoux, 2021). The white worker’s bargaining power, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012). Consistent with our expectations, we find that the black workers have a lower bargaining power compared to the white workers, with $\xi_{bl}$ to be 0.1397 and $\xi_{wh}$ to be 0.2110.
@@ 3.6 Assessing the model as a quantitative theory of racial disparity
Our calibration successfully captures the racial gaps across three important dimensions: income, wealth, and employment outcomes. We now discuss these dimensions in more detail before proceeding with the quantitative exercises that examine the macroeconomic impact of racial discrimination.
@@
Figure 1: Labor Income and Wealth Distribution
Note: This figure compares the model-generated steady-state distributions of wealth and labor income with their empirical analogs. The empirical labor income and wealth distributions are estimated from the 2007 American Community Survey (ACS) and the 2007 Consumer Finance Survey (SCF).
First, the model captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we do not target these distributions explicitly. As with most incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile. On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
@@
Mean wealth ratio 0.23 0.28
@@
Note: This table compares the steady state wealth and unemployment moments between black and white workers with the corresponding empirical data.
Second, the model reproduces the disparities in unemployment, labor income, and wealth between black and white workers. As targeted in the calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth. In addition to the targeted moments, Table 2 demonstrates that our model successfully generates the untargeted employment and wealth moments. Specifically, our model produces a 5% unemployment rate among white workers compared to 12% among black workers. Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to prejudiced firms. On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.28, close to 0.23 as in Kaplan et al. (2014). However, the model understates the median wealth gap between black and white workers. The median black worker holds 33% of the wealth of the median white worker in our model compared to 17% in Kaplan et al. (2014).
In summary, the model reflects the racial disparities present in labor income, wealth, and unemployment outcomes. Next, we will investigate how racial discrimination in hiring contributes to these disparities.
@@ 4 Steady state results
In this section, we first establish how a search framework sustains hiring discrimination in equilibrium. Then, we examine the steady-state impact of racial discrimination by comparing the benchmark economy to an alternative economy without hiring discrimination. Lastly, we explore the heterogeneous welfare implications associated with eliminating hiring discrimination.
@@ 4.1 Sustaining racial discrimination
The labor search-and-matching framework allows firms to earn a profit from the surplus generated by search frictions. This feature permits prejudiced and non-prejudiced firms to coexist in a competitive equilibrium. It contrasts with a canonical frictionless neoclassical model where prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers as described in Becker (1957).
Table 3 shows firm performances in the steady state. $p$ firms have a total profit of 0.04, while $np$ firms have 0.03. This difference arises because $p$ firms experience a lower job destruction shock (0.0268) than $np$ firms (0.0644). Although $p$ firms incur higher vacancy posting costs (4.2622) to hire only white workers, they retain workers longer (employment level at 0.54) at lower posting rates (0.02). Conversely, $np$ firms face lower vacancy posting costs (2.735), but their high job destruction shock leads to a lower employment level (0.40) and a more frequent need for vacancy postings (0.05).
@@
p firm employment 0.54
np firm employment 0.40
Note: This table shows steady state profit, vacancy postings, and employment levels for $p$ and $np$ firms in the benchmark model.
Next, we examine the steady-state interplay between prejudiced and non-prejudiced firms. At the benchmark equilibrium, $p$ firms post a positive number of vacancies with sufficient expected profits to cover the posting costs. We gradually increase penalties for them by raising $\kappa_{p}$ until their expected operating profit no longer justifies vacancy postings. Figure 2 displays the profit, vacancy, and employment dynamics between the $p$ and $np$ firms. As $\kappa_{p}$ rises, $p$ firms become increasingly disadvantaged, gradually losing their profits, vacancy postings, and employment. Conversely, with fewer $p$ firms competing in the market, $np$ firms see a steady increase in profits, vacancy postings, and employment.
It is worth noting that the dynamics of $p$ and $np$ firms are not symmetric, given the differences in vacancy posting costs and job destruction shocks. As $p$ firms’ profits decline, the profits of $np$ firms increase at a slower rate. Additionally, as $np$ firms replace $p$ firms, $np$ firms recruit from both black and white workers, rather than only white workers displaced from $p$ firms. This is mainly reflected in $np$ firms’ more elastic vacancy posting response.
In the limit, as $\kappa_{p}$ approaches the total expected value of an operating $p$ firm, posting vacancies exclusively for white workers becomes entirely unprofitable. As a result, all $p$ firms exit the market. The value of $\kappa_{p}$ at which $p$ firms exit entirely is 217% of the benchmark value of 4.2622. Meanwhile, $np$ firms become the sole employers in the economy. The gray dotted lines in Figure 2 illustrate a steady state where we construct a counterfactual model that structurally removes $p$ firms from the benchmark model. Thus, increasing $\kappa_{p}$ in the limit asymptotically equals the alternative of structurally eliminating $p$ firms.
Figure 2: Penalizing discrimination
Note: This figure displays the dynamics between $p$ and $np$ firms as $\kappa_{p}$ increases. The top panel plots the change in profit, the middle panel plots the change in vacancy posting, and the bottom panel plots the change in employment. The red dotted line depicts the dynamics of $p$ firms, the blue solid line depicts the dynamics of $np$ firms, and the gray dotted line depicts the limit where no $p$ firm exists. The horizontal axes record the values of $\kappa_{p}$ from the benchmark $\kappa_{p}$ at 4.2622 to 217% of the benchmark value, at 9.2444. The vertical axes record the model-generated value of each variable.
@@ 4.2 Impact of hiring discrimination on racial inequality
In this subsection, we compare the benchmark economy to those without prejudiced firms to illustrate the aggregate and distributional impact of discriminatory hiring. Table 4 presents the comparisons. All firms in the counterfactual models provide equal hiring opportunities to black and white workers. Column 2 removes the $p$ firms from the benchmark model, which leaves all workers with equal job-finding rates with the $np$ firms and equal job destruction shocks at the benchmark $\lambda_{np}$. In Column 3, we further adjust the job destruction shocks to be race-specific, with $\lambda_{wh}$ and $\lambda_{bl}$ designating the shocks for white and black workers, respectively. [^8] In the benchmark economy, we structure firm-specific job destruction shocks ($\lambda_{np}$ and $\lambda_{p}$) to match race-specific job separation rates. This counterfactual economy only has one type of firm. We assign white workers a job destruction shock, $\lambda_{wh}$, of 0.0380 and black workers a destruction shock, $\lambda_{bl}$, of 0.0644 to maintain black and white workers’ job separation rates as in the calibration. Reassigning the job destruction shocks further isolates the impact of job separation.
@@
Moments Benchmark Eliminate $p$ firms Race specific $\lambda$
@@
job separation rate - black 0.06 0.06 0.06
job separation rate - white 0.04 0.06 0.04
job finding rate - black 0.50 0.59 0.62
job finding rate - white 0.66 0.59 0.62
unemp rate - black 0.12 0.10 0.09
unemp rate - white 0.05 0.10 0.06
mean wage - black 1.60 1.96 1.82
mean wage - white 2.13 2.22 2.11
mean wage ratio 0.75 0.88 0.86
mean wealth ratio 0.28 0.33 0.33
median wealth ratio 0.33 0.44 0.44
@@
$p$ firm profit 0.04 - -
np firm profit 0.03 0.06 0.06
p firm vacancy 0.02 - -
np firm vacancy 0.05 0.10 0.07
p firm employment 0.54 - -
np firm employment 0.40 0.90 0.94
@@
$p$ market tightness 0.40 - -
$np$ market tightness 0.73 1.02 1.12
@@
Y 3.06 2.95 3.06
K/Y 10.26 10.22 10.25
average wage 2.04 1.96 2.05
unemp rate 0.07 0.10 0.06
@@
Note: This table compares the benchmark steady state with a counterfactual steady state without $p$ firms. In the counterfactual model, all firms provide equal hiring opportunities to black and white workers and keep race-specific job destruction shock. All other parameters stay the same as benchmark calibration.
The first panel illustrates the changes in racial employment, wage, and wealth disparities after removing prejudiced firms. The first two rows describe the job separation rates. By construction, Column 2 equalizes job separation rates by race, while Column 3 retains the differences as the benchmark calibration. The next two rows describe the job-finding rates. Since there are no $p$ firms in the economy in Columns 2 and 3, the job-finding rates are equal for black and white workers. White workers have only one sector for employment; hence, their job-finding rate decreases. Black workers have more job opportunities as the $np$ sector posts more vacancies; thus, their job-finding rate increases. Comparing Columns 2 and 3, the much lower job destruction shock for white workers in Column 3 results in fewer workers searching for jobs, which leads to higher job-finding rates than in Column 2. Altogether, removing $p$ firms reduces the racial gap in the unemployment rate. In Column 2, the racial unemployment rate gap completely closes as job turnover equalizes. In Column 3, the unemployment rate gap closes from 7% to 3% due to the remaining racial difference in $\lambda$.
@@
Turning to the wage impact, black workers can bargain for higher wage rates as they receive more job opportunities when we remove discriminatory hiring. White workers, however, no longer have access to the exclusive $p$ sector; their outside options are lower during bargaining, thus driving down their wage rates. As a result, the average black-to-white wage ratio increases from 75% in the benchmark to 88% in Column 2 and to 86% in Column 3. We conclude that removing discriminatory hiring closes 52% (Column 2) to 44% (Column 3) of the average wage gap between black and white workers.
Additionally, removing discriminatory hiring reduces the wealth gap, with the mean wealth ratio moving from 0.28 to 0.33 and the median wealth ratio increasing from 0.33 to 0.44. The results remain stable between Columns 2 and 3. We infer that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap.
The remaining gaps in wage and wealth come from the bargaining power and extreme wealth shock disparities between black and white workers. [^9] Though not the primary purpose of this paper, we present the results after removing racial differences in bargaining power and extreme wealth shock in Appendix B.
The second panel displays the changes in firm attributes. [^10] We loosely refer to the one-sector economy without discriminatory hiring as $np$ firms, though the Column 3 specification differs from the benchmark $np$ firms in their job destruction shock. Without prejudiced firms, $np$ firms’ profits double, and vacancy postings and employment levels increase. These changes result in the aforementioned advantages for black workers in the counterfactual economies. In Column 3, the lower job destruction shock for white workers further drives up the overall $np$ employment and reduces vacancy postings compared to Column 2.
The third panel of Table 4 displays the labor market conditions. As $p$ firms exit, a greater number of $np$ firms enter the market endogenously. The labor market tightness for the $np$ sector, $\theta_{np}$, jumps from 0.73 in the benchmark to 1.02 in the equilibrium without $p$ firms, and to 1.12 when we further split $\lambda$ by race. The ensuing market tightness is higher than the benchmark aggregate market tightness of 1.
The last panel of Table 4 describes the aggregate outcomes. With more $np$ firms entering, the aggregate output remains relatively unchanged in Column 3. The capital-output ratio slightly decreases to 10.25. The economy-wide average wage increases to 2.05 from the benchmark of 2.04, and the aggregate unemployment rate decreases to 6%. However, Column 2 experiences a more significant reduction in output and average wage, along with a higher unemployment rate due to the overall higher job destruction shock that requires firms to frequently pay for vacancy posting costs instead of using resources for production.
Our results show that the presence of prejudiced firms can explain sizable fractions of income and wealth inequality. Penalizing prejudiced firms primarily improves employment opportunities and wage outcomes for black workers; it also worsens the situation for white workers. Removing $p$ firms decreases job-finding rates and increases unemployment for white workers, lowering their bargaining positions and affecting their income and wealth. To quantify the differential effects of removing discriminatory hiring, we next calculate the welfare changes experienced by black and white workers.

### v3

@@ 3.1 Preferences
We set the period utility function $u(c)$ to be $\log(c)$. The discount factor, $\beta$, is calibrated to match the quarterly capital-output ratio of 10.26, a value used by a number of studies such as Den Haan et al. (2010) and Carroll et al. (2017). The resulting value of $\beta$ is 0.9943, and the corresponding quarterly real interest rate is 1.3%.
@@ 3.2 Production
The matched worker-firm pairs produce according to a Cobb-Douglas production function, $f(k)=k^{\alpha_{k}}$. We choose $\alpha_{k}$ to be 0.289 and set the quarterly capital depreciation rate $\delta$ equal to 0.015, following Nakajima (2012).
@@ 3.3 Productivity and wealth shocks
We use hourly real wages from the PSID to estimate the persistence, $\rho_{s}$, and the standard deviation, $\sigma_{s}$ of the productivity process. Our estimation strategy follows Setty and Yedid-Levi (2021). We run a Mincer regression of log wages on demographic controls and year and state fixed effects. The demographic controls include education, labor market experience, race, and marital status. We restrict the sample to men aged 25 and older. We then use the residuals obtained to estimate an AR(1) regression using the panel dimension of the PSID data. The estimated regression operates on a biennial frequency since the PSID data is released every two years. We assign the quarterly adjusted coefficient of the AR(1) regression as the persistence of the idiosyncratic productivity process. The standard deviation of the productivity process corresponds to the standard deviation of the residuals from the AR(1) regression after adjusting for the model frequency. We estimate the quarterly persistence, $\rho_{s}$, to be 0.9411 and the corresponding standard deviation, $\sigma_{s}$, to be 0.1680.666We follow Setty and Yedid-Levi (2021) to have only men in the sample. We also estimate the AR(1) process to include both men and women from PSID. The resulting $\rho_{s}$ is unchanged (0.9411), and $\sigma_{s}$ is slightly lower (0.1671) than the men-only sample.
Following Mukoyama (2010), our model also features extreme wealth shocks to capture the mass of people with zero wealth. The race-specific probability of losing one’s wealth, $\epsilon_{R}$, is calibrated to be 0.0179 for black workers and 0.0088 for white workers. They capture the empirical moments that around 18% of black workers and 7% of white workers have zero wealth (Nakajima, 2021).
@@ 3.4 Unemployment insurance
The unemployment insurance system in our model is characterized by the replacement rate, $h$, the maximum insurance payout, $\chi$, and the probability of maintaining eligibility status, $P_{e}$. We choose $h$ to be 0.4 following Shimer (2005) and Mitman and Rabinovich (2015). In accordance with Setty and Yedid-Levi (2021), we calibrate $\chi$ to 0.8403, which amounts to 48% of the median wage in the model. The eligibility probability $P_{e}$ is set to 0.5385 to generate an average duration of unemployment benefits of 26 weeks, as in Mitman and Rabinovich (2015).
@@ 3.5 Labor search and wage bargain
The parameters governing the labor search are calibrated to match the labor market turnover statistics obtained from Cajner et al. (2017). The elasticity of the matching function, $\iota$, targets the job-finding rate of black workers. Our calibrated value of 1.3452 is close to 1.25, the value used by Den Haan et al. (2000) and Petrosky-Nadeau et al. (2018). We choose firm-type specific job destruction rates $\lambda_{np}$ and $\lambda_{p}$ to match the job separation rates of black and white workers, respectively. The ensuing values for $\lambda_{np}$ and $\lambda_{p}$ are 0.0644 and 0.0268, respectively. [^7] All the employed black workers only work with $np$ firms, and hence $\lambda_{np}$ is exactly equal to the job separation rate of black workers. White workers are employed in both $p$ and $np$ firms. Hence, $\lambda_{p}$ (0.0268) is lower than the aggregate separation rate of white workers (0.0380). The vacancy posting cost of prejudiced firms, $\kappa_{p}$, is chosen to match the job-finding rate of white workers. The posting cost of non-prejudiced firms, $\kappa_{np}$, is chosen to match the aggregate labor market tightness $\theta$, following Wolcott (2021). [^8] Aggregate labor market tightness is defined as $\theta=\frac{v_{np}+v_{p}}{u_{bl}+u_{wh}}$. We find that prejudiced firms pay significantly more than non-prejudiced firms to post their vacancies, with $\kappa_{p}$ calibrated at 4.6314, compared to $\kappa_{np}$ at 2.5534. The bargaining power of black workers, $\xi_{bl}$, is calibrated to target the average black-white racial wage ratio of 0.75 (Derenoncourt and Montialoux, 2021). The bargaining power of the white worker, $\xi_{wh}$, is chosen to generate the average firm profit share of 3.3% as in Nakajima (2012). Consistent with our expectations, black workers have a lower bargaining power than white workers, with $\xi_{bl}$ calibrated to be 0.1371 and $\xi_{wh}$ to be 0.1988.
@@ 3.6 Assessing the model as a quantitative theory of racial disparity
Our calibration successfully captures the racial gaps across three important dimensions: income, wealth, and employment outcomes. We now discuss the model validity in more detail before proceeding with the quantitative exercises that examine the macroeconomic impact of racial discrimination.
@@
Figure 1: Labor income and wealth distributions
Note: This figure compares the steady-state distributions of wealth and labor income generated by the model with their empirical counterparts. The empirical labor income and wealth distributions are estimated using data from the 2007 American Community Survey (ACS) and the 2007 Consumer Finance Survey (SCF), respectively. The horizontal axes indicate quintiles, from the lowest (Q1) to the highest (Q5), while the vertical axes report the share of total income or wealth accounted for by each quintile. Blue bars represent the data, and red dotted bars show the model-simulated moments.
First, the model successfully captures the overall household distribution of wealth and labor earnings as shown in Figure 1, even though we do not explicitly target these distributions. As with most incomplete market models, we also face difficulty in generating the extreme concentration of wealth and labor income in the top quintile of the distribution. On the other hand, we are much closer to the empirical distributions in the lower quintiles since we target the share of zero-wealth workers in our calibration.
@@
Mean wealth ratio 0.23 0.29
@@
Note: This table compares the empirical unemployment and wealth moments with the corresponding steady-state moments generated by the model.
Second, the model reproduces the disparities in unemployment, labor income, and wealth between black and white workers. Through calibration, the model replicates the empirical racial labor income gap and the empirical racial difference in the share of people at zero wealth. In addition to the targeted moments in calibration, Table 2 shows that our model successfully generates the untargeted employment and wealth moments close to the data. Specifically, our model produces a 5% unemployment rate among white workers compared to 12% among black workers. Apart from the lower separation rate, white workers have a lower unemployment rate due to their access to prejudiced firms. On the wealth dimension, without targeting, the model generates a mean black-white wealth ratio of 0.29, close to 0.23 in the data (Kaplan et al., 2014). However, the model underestimates the median wealth gap between black and white workers. The median black worker holds 33% of the wealth of the median white worker in our model compared to the empirical value of 17% (Kaplan et al., 2014).
In summary, the model reflects the racial disparities present in labor income, wealth, and unemployment outcomes. Next, we investigate how racial discrimination in hiring contributes to these disparities.
@@ 4 Steady state results
In this section, we first establish how a search framework sustains hiring discrimination in equilibrium. Then, we examine the steady-state impact of racial discrimination by comparing the benchmark economy with an alternative economy without hiring discrimination. Lastly, we explore the heterogeneous welfare implications associated with eliminating hiring discrimination.
@@ 4.1 Sustaining racial discrimination
The presence of search-and-matching frictions allows firms to earn a profit from the surplus generated by search frictions. This feature permits prejudiced and non-prejudiced firms to coexist in the equilibrium. This is in contrast to a canonical frictionless neoclassical model in which prejudiced hiring will be driven out by competition if there is no inherent difference between black and white workers, as discussed in Becker (1957).
@@
p firm employment 0.57
np firm employment 0.37
Note: This table presents steady state profit, vacancy postings, and employment levels for $p$ and $np$ firms in the benchmark model.
Table 3 shows the performance of firms in the steady state. Prejudiced firms earn a total profit of 0.04, while the corresponding profit of non-prejudiced firms is 0.02. This difference arises because $p$ firms experience a lower job destruction shock (0.0268) compared to $np$ firms (0.0644). Even though $p$ firms incur higher vacancy posting costs (4.6314) to hire only white workers, they retain their workers for a longer period, reducing the number of vacancy postings. In contrast, $np$ firms face lower vacancy posting costs (2.5534), but their higher job destruction rate leads to a more frequent need for vacancy postings (0.05). In the steady state, $p$ firms sustain an employment level (0.57) higher than $np$ firms (0.37).
@@ 4.2 Impact of hiring discrimination on racial inequality
In this subsection, we analyze the distributional and aggregate effects of discriminatory hiring by comparing the benchmark economy to a counterfactual economy without prejudiced firms. The counterfactual economy contains only non-prejudiced firms, where both black and white workers retain their race-specific job separation rates from the benchmark calibration. By holding these separation rates constant, we isolate how discrimination in hiring - rather than differences in job separation - shapes labor market outcomes. [^9] The benchmark economy calibrates prejudiced and non-prejudiced firms’ job destruction shocks ($\lambda_{p}$ and $\lambda_{np}$) to match observed racial disparities in job separation rates. In the single-firm-type counterfactual, we preserve benchmark separation rates by assigning white workers a $\lambda_{wh}$ of 0.0380 and black workers a $\lambda_{bl}$ of 0.0644. Table 4 presents the comparison.
@@
Moments Benchmark No $p$ firms
@@
job separation rate - black 0.06 0.06
job separation rate - white 0.04 0.04
job finding rate - black 0.49 0.64
job finding rate - white 0.66 0.64
unemp rate - black 0.12 0.09
unemp rate - white 0.05 0.06
mean wage - black 1.59 1.85
mean wage - white 2.12 2.11
mean wage ratio 0.75 0.87
mean wealth ratio 0.29 0.34
median wealth ratio 0.33 0.44
@@
$p$ firm profit 0.04 -
np firm profit 0.02 0.06
$p$ firm vacancy 0.02 -
np firm vacancy 0.05 0.07
$p$ firm employment 0.57 -
np firm employment 0.37 0.94
@@
$p$ market tightness 0.39 -
$np$ market tightness 0.71 1.17
@@
Y 3.07 3.08
K/Y 10.26 10.27
average wage 2.04 2.07
unemp rate 0.06 0.06
@@
Note: This table compares the benchmark steady state with a counterfactual steady state without $p$ firms. In the counterfactual model, all firms provide equal hiring opportunities to black and white workers. All other parameters stay the same as benchmark calibration.
The first panel shows how racial disparities in household outcomes change when prejudiced firms are removed. While the racial differences in job separation rates remain unchanged at benchmark levels, eliminating discriminatory hiring practices leads to equalized job-finding rates for black and white workers. White workers experience a slight decrease in their job-finding rate, dropping from 0.66 to 0.64, due to losing access to $p$ firms. Meanwhile, black workers’ job-finding rate rises significantly, from 0.49 to 0.64, in the counterfactual equilibrium. The removal of $p$ firms leads to endogenous entry of $np$ firms, which increases the number of vacancies and thus raises the job-finding rate for black workers. Overall, this removal in prejudiced firms narrows the racial unemployment gap from 7 percentage points to 3 percentage points, showing that discriminatory hiring explains approximately 57% of the unemployment gap between black and white workers.
@@
Turning to the impact on wages, removing hiring discrimination boosts outcomes for black workers by improving their outside options. This enables black workers to bargain for higher wages upon a successful match. However, white workers no longer have access to jobs from the exclusive $p$ firms. This lowers their outside option during bargaining, slightly reducing their wage rates. As a result, the average wage ratio between black and white workers increases from 0.75 to 0.87. We can conclude that hiring discrimination accounts for roughly 48% of the wage gap.
The change in hiring practices also narrows the wealth gap: the mean wealth ratio rises from 0.29 to 0.33, and the median wealth ratio increases from 0.34 to 0.44. These shifts imply that discriminatory hiring explains about 7% of the mean wealth gap and 16% of the median wealth gap. The remaining differences in wages and wealth can be accounted for by differences in bargaining power and extreme wealth shocks between black and white workers. [^10] Although not the primary focus of this paper, additional results that remove racial differences in bargaining power and extreme wealth shocks are presented in Appendix B.
The second panel examines firm dynamics when discriminatory firms are removed. [^11] We refer to the one-sector economy without discriminatory hiring as $np$ firms, although their job destruction rate $\lambda$ differs from that of the benchmark $np$ firms. Without prejudiced firms, $np$ firms see their profits triple, which encourages additional $np$ firms to enter the market and post vacancies. Consequently, $np$ firms fully replace $p$ firms in the counterfactual equilibrium. As shown in Table 4, the total number of vacancies and overall employment remain unchanged compared to the benchmark. Still, eliminating prejudiced firms helps narrow the racial gap in labor market outcomes, as described earlier.
The third panel of Table 4 shows how labor market conditions adjust when $p$ firms exit. As these firms leave, additional $np$ firms enter endogenously, raising the labor market tightness in the $np$ market from 0.71 in the benchmark to 1.17 in the counterfactual equilibrium. Consequently, the overall market tightness exceeds the benchmark aggregate level of 1.
The last panel of Table 4 presents the aggregate outcomes. Despite the increased entry of $np$ firms, aggregate output rises only marginally from 3.07 to 3.08. The capital-output ratio increases from 10.26 to 10.27, and the aggregate wage grows from 2.04 to 2.07, while the aggregate unemployment rate remains steady at 6%.
Our results suggest that prejudiced firms account for a substantial share of employment, income, and wealth inequality. Eliminating $p$ firms improves employment and wage outcomes for black workers, while white workers experience a slight decline in their job-finding rate and an increase in unemployment. This reduction in job prospects weakens white workers’ bargaining power and adversely affects their income and wealth. In the next section, we quantify these differential effects by calculating the welfare changes experienced by black and white workers.